# Uniqueness of invariant measures as a structural property of Markov kernels

### v1

@@ Abstract
Once existence of invariant probability measures is known, indecomposability alone is sufficient to rule out multiplicity. On standard Borel spaces, this viewpoint is consistent with the classical theory: irreducibility appears as a convenient sufficient condition ensuring indecomposability, rather than as a structural requirement for uniqueness.
@@ 2 Main result
We first recall a key structural property of invariant ergodic measures. Although the following result can be derived from the ergodic decomposition theorem under standard assumptions, we include a direct proof since our goal is to isolate the measure-theoretic core of the uniqueness mechanism, without relying on convex-analytic or topological arguments on the space of invariant measures.
@@
Let $(E,\mathcal{B})$ be a standard Borel space and let $P$ be a Markov kernel on $(E,\mathcal{B})$. If $P$ admits two distinct invariant probability measures, then it admits two invariant probability measures that are mutually singular.
@@
Assume that $P$ admits two distinct invariant probability measures. Let $\mu_{1}$ and $\mu_{2}$ be two such measures with $\mu_{1}\neq\mu_{2}$. Then there exists a measurable set $A\in\mathcal{B}$ such that $\mu_{1}(A)\neq\mu_{2}(A)$.
By the ergodic decomposition theorem, every invariant probability measure admits a representation as a barycenter of invariant ergodic probability measures. Since $\mu_{1}\neq\mu_{2}$, their ergodic decompositions differ, and there exist two distinct invariant ergodic probability measures $\nu_{1}$ and $\nu_{2}$ such that $\nu_{1}$ appears with positive weight in the decomposition of $\mu_{1}$ and $\nu_{2}$ appears with positive weight in the decomposition of $\mu_{2}$. Since distinct invariant ergodic probability measures are mutually singular, this yields two invariant probability measures that are mutually singular.
@@
The singularity of distinct invariant ergodic probability measures, as well as the extremality of ergodic measures, are classical results, often derived from the Choquet simplex structure of the set of invariant probability measures; see, e.g., [10, 9]. The argument above relies only on standard measure–theoretic properties of invariant measures and does not involve any trajectorywise ergodic theorem.
Lemma 3 shows that invariant ergodic measures are necessarily extreme points of the convex set of invariant probability measures. We now record a complementary observation, showing that ergodicity is automatic under uniqueness.
@@
Since $P\mathbf{1}_{A}=\mathbf{1}_{A}$ $\mu$-a.s. and $0<\mu(A)<1$, the measures $\mu_{1}$ and $\mu_{2}$ are well defined invariant probability measures, with disjoint supports. In particular, $\mu_{1}\neq\mu_{2}$, which contradicts the uniqueness assumption. Therefore $\mu(A)\in\{0,1\}$ for every $A\in\mathcal{B}$ such that $P\mathbf{1}_{A}=\mathbf{1}_{A}$ $\mu$-almost surely, and $\mu$ is ergodic. ∎
@@
We now turn to the complementary situation where several invariant probability measures coexist.
Corollary 7.
Let $(E,\mathcal{E})$ be a measurable space and let $P$ be a Markov kernel on $(E,\mathcal{E})$. If $P$ admits at least two distinct invariant probability measures, then it admits two distinct invariant probability measures that are mutually singular.
Proof.
Assume that $P$ admits two distinct invariant probability measures $\mu_{1}\neq\mu_{2}$.
If both $\mu_{1}$ and $\mu_{2}$ are ergodic, then by Lemma 3 they are mutually singular and the conclusion holds.
Otherwise, at least one invariant probability measure is not ergodic. Without loss of generality, assume that $\mu_{1}$ is not ergodic. Then there exists a measurable set $A\in\mathcal{E}$ such that
$P\mathbf{1}_{A}=\mathbf{1}_{A}\quad\mu_{1}\text{-a.s.}\qquad\text{and}\qquad 0<\mu_{1}(A)<1.$
Define the probability measures
$\nu_{1}(B):=\frac{\mu_{1}(B\cap A)}{\mu_{1}(A)},\qquad\nu_{2}(B):=\frac{\mu_{1}(B\cap A^{c})}{1-\mu_{1}(A)},\qquad B\in\mathcal{E}.$
Since $P\mathbf{1}_{A}=\mathbf{1}_{A}$ $\mu_{1}$-almost surely, the measures $\nu_{1}$ and $\nu_{2}$ are invariant probability measures for $P$. Moreover, they have disjoint supports and are therefore mutually singular. ∎
Remark 8.
The proof relies only on absolute continuity and invariant sets, together with standard structural properties of invariant measures. Ergodicity appears here as extremality within a fixed domination class.
@@
Let $P$ be a Markov kernel on a standard Borel space $(E,\mathcal{B})$. If $P$ is indecomposable, then $P$ admits at most one invariant probability measure.
@@
Assume by contradiction that $P$ admits two distinct invariant probability measures. By Corollary 7, there exist two mutually singular invariant probability measures $\mu_{1}$ and $\mu_{2}$. Hence there exists $A\in\mathcal{B}(E)$ such that
@@
Thus $B_{1}$ and $B_{2}$ are disjoint nontrivial absorbing measurable sets, contradicting the indecomposability of $P$. ∎
@@
On a standard Borel space, $\phi$–irreducibility in the sense of Meyn–Tweedie is a strictly stronger property than indecomposability, since $\phi$–irreducibility rules out the existence of nontrivial absorbing measurable sets. In general, the converse implication does not hold, as indecomposability alone does not preclude purely transient behavior. However, if an invariant probability measure exists, the structural decomposition theory of [10, Section 4.2] implies that indecomposability excludes the presence of more than one closed communicating class and therefore enforces irreducibility with respect to a maximal irreducibility measure $\psi$ in the sense of Meyn–Tweedie. In this sense, indecomposability and $\phi$–irreducibility become equivalent once the existence of an invariant probability measure is guaranteed. From this perspective, indecomposability identifies a purely structural obstruction to uniqueness of invariant probability measures, while stronger assumptions—such as the existence of small or petite sets or additional regularity of the transition kernel—are only required to address recurrence and ergodic convergence properties.
@@ 2.3 Stability via uniqueness
Theorem 17 (Stability under uniqueness).
Assume that:
  1. 1.
    the transition kernel $P$ is quasi–Feller;
  2. 2.
    the empirical occupation measures $(\Lambda_{n})$ are $P_{x}$–almost surely tight for every $x\in E$;
  3. 3.
    $P$ admits a unique invariant probability measure $\mu$.
Then the Markov chain $(X_{n})_{n\geq 0}$ is stable in the sense of Duflo.
@@ 4 Conclusion
We further show that, under irreducibility, indecomposability can be enforced through a resolvent–type kernel. The resolvent transforms finite–time reachability into a one–step positivity property at the level of invariant measures, forcing all invariant probability measures to charge a common $\sigma$–finite reference measure. Since distinct invariant ergodic measures are mutually singular, this mechanism excludes the possibility of multiple ergodic components and yields uniqueness without any appeal to return–time estimates or petite set constructions.

### v2

@@ Abstract
Once existence of invariant probability measures is established, indecomposability alone is sufficient to rule out multiplicity. On standard Borel spaces, this viewpoint is consistent with the classical theory: irreducibility appears as a convenient sufficient condition ensuring indecomposability, rather than as a structural requirement for uniqueness.
@@ 2 Main result
We first recall a fundamental structural property of invariant ergodic measures. Although the following result can be derived from the ergodic decomposition theorem under standard assumptions, we include a direct proof since our goal is to isolate the measure–theoretic core of the uniqueness mechanism, without relying on convex–analytic or topological arguments on the space of invariant measures. In particular, an observation underlying all subsequent arguments is that distinct invariant ergodic probability measures must be mutually singular.
@@
Let $P$ be a Markov kernel on a measurable space $(E,\mathcal{E})$. Any two distinct PP–invariant ergodic probability measures are mutually singular.
@@
Let $\mu$ and $\nu$ be two distinct PP–invariant ergodic probability measures and set
$\eta:=\tfrac{1}{2}(\mu+\nu).$
Then $\eta$ is PP–invariant and $\mu\ll\eta$. Let $f:=d\mu/d\eta$. Since $\mu+\nu=2\eta$, we also have $d\nu/d\eta=2-f$, hence $0\leq f\leq 2$ $\eta$–a.s.
Step 1: $Pf=f$ $\eta$–a.s. Fix a bounded measurable function $\varphi\in L^{\infty}(E)$. Using $\mu=f\,\eta$ and the invariance of $\mu$, we get
$\int_{E}\varphi(x)\,f(x)\,\eta(dx)=\int_{E}\varphi(x)\,\mu(dx)=\int_{E}(P\varphi)(x)\,\mu(dx)=\int_{E}(P\varphi)(x)\,f(x)\,\eta(dx).$ (1)
We now rewrite the last term. Define the finite measure $\Pi$ on $(E\times E,\mathcal{E}\otimes\mathcal{E})$ by
$\Pi(dx,dy):=\eta(dx)\,P(x,dy).$
Since $\varphi$ is bounded and $f\in L^{1}(\eta)$ with $f\geq 0$, the function
$(x,y)\longmapsto|\varphi(y)|\,f(x)$
is $\Pi$–integrable:
$\int_{E\times E}|\varphi(y)|\,f(x)\,\Pi(dx,dy)=\int_{E}f(x)\left(\int_{E}|\varphi(y)|\,P(x,dy)\right)\eta(dx)\leq\|\varphi\|_{\infty}\int_{E}f\,d\eta<\infty.$
Hence Tonelli’s theorem applies and yields
$\int_{E}(P\varphi)(x)\,f(x)\,\eta(dx)$ $=\int_{E}\left(\int_{E}\varphi(y)\,P(x,dy)\right)f(x)\,\eta(dx)$
$=\int_{E\times E}\varphi(y)\,f(x)\,\Pi(dx,dy)$
$=\int_{E}\varphi(y)\left(\int_{E}f(x)\,P(x,dy)\,\eta(dx)\right).$
By definition of the action of $P$ on $L^{1}(\eta)$ (equivalently, by the duality $\int(P\varphi)\,f\,d\eta=\int\varphi\,Pf\,d\eta$), the last expression equals $\int_{E}\varphi(y)\,(Pf)(y)\,\eta(dy)$. Plugging this into (1) gives
$\int_{E}\varphi\,f\,d\eta=\int_{E}\varphi\,Pf\,d\eta\qquad\forall\varphi\in L^{\infty}(E),$
hence $Pf=f$ $\eta$–a.s.
Step 2: $f$ is constant $\mu$–a.s. and $\nu$–a.s. Since $Pf=f$ $\mu$–a.s. and $0\leq f\leq 2$, ergodicity of $\mu$ implies that $f$ is $\mu$–a.s. constant: $f=c$ $\mu$–a.s. for some $c\in[0,2]$. Similarly, $f=c^{\prime}$ $\nu$–a.s. for some $c^{\prime}\in[0,2]$.
If $c=c^{\prime}$, then $f$ is $\eta$–a.s. constant, hence $\mu=f\eta$ and $\nu=(2-f)\eta$ are both proportional to $\eta$; since they are probability measures, this forces $\mu=\nu$, a contradiction. Therefore $c\neq c^{\prime}$.
Step 3: the only possible constants are $c=2$ and $c^{\prime}=0$. Since $\mu$ is supported on $\{f=c\}$, we have
$1=\mu(E)=\int_{\{f=c\}}f\,d\eta=c\,\eta(f=c),\qquad\text{so}\qquad\eta(f=c)=\frac{1}{c}.$
Likewise, $\nu$ is supported on $\{f=c^{\prime}\}$ and $d\nu=(2-f)\,d\eta$, hence
$1=\nu(E)=\int_{\{f=c^{\prime}\}}(2-f)\,d\eta=(2-c^{\prime})\,\eta(f=c^{\prime}),\qquad\text{so}\qquad\eta(f=c^{\prime})=\frac{1}{2-c^{\prime}}.$
The sets $\{f=c\}$ and $\{f=c^{\prime}\}$ are disjoint (since $c\neq c^{\prime}$), thus
$1=\eta(E)\geq\eta(f=c)+\eta(f=c^{\prime})=\frac{1}{c}+\frac{1}{2-c^{\prime}}.$
If $c<2$ then $1/c>1/2$, and if $c^{\prime}>0$ then $1/(2-c^{\prime})>1/2$, so the right-hand side would be >1>1, a contradiction. Therefore $c=2$ and $c^{\prime}=0$.
Step 4: conclusion. Let $A:=\{f=2\}$. Then $\mu(A)=1$ and $\nu(A)=0$, hence $\mu\perp\nu$. ∎
@@
Although the statement of Lemma 3 is reminiscent of classical results based on ergodic decompositions and Choquet theory, the proof given here is entirely independent of such tools. It relies only on invariance, absolute continuity, and the fact that bounded $P$-harmonic functions are constant under ergodicity. In particular, no topological or convex-analytic structure on the set of invariant measures is required.
Proposition 5.
If a Markov kernel $P$ admits at least two distinct invariant probability measures, then it admits two invariant probability measures that are mutually singular.
Proof.
Let $\mu\neq\nu$ be two invariant probability measures.
If both $\mu$ and $\nu$ are ergodic, the result follows directly from Lemma 3.
Otherwise, at least one of them, say $\mu$, is not ergodic. Then there exists a measurable set $A\in\mathcal{E}$ such that
$0<\mu(A)<1\quad\text{and}\quad P(x,A)=\mathbf{1}_{A}(x)\quad\mu\text{-a.s.}$
Define the conditional probability measures
$\mu_{A}(B):=\frac{\mu(B\cap A)}{\mu(A)},\qquad\mu_{A^{c}}(B):=\frac{\mu(B\cap A^{c})}{\mu(A^{c})}.$
Since $A$ is $\mu$-almost surely absorbing, both $\mu_{A}$ and $\mu_{A^{c}}$ are invariant probability measures. Moreover, they are mutually singular since
$\mu_{A}(A)=1\quad\text{and}\quad\mu_{A^{c}}(A)=0.$
This concludes the proof. ∎
@@
Since $P\mathbf{1}_{A}=\mathbf{1}_{A}$ $\mu$-a.s. and $0<\mu(A)<1$, the measures $\mu_{1}$ and $\mu_{2}$ are well defined invariant probability measures and are mutually singular (indeed $\mu_{1}(A)=1$ and $\mu_{2}(A)=0$). In particular, $\mu_{1}\neq\mu_{2}$, which contradicts the uniqueness assumption. Therefore $\mu(A)\in\{0,1\}$ for every $A\in\mathcal{B}$ such that $P\mathbf{1}_{A}=\mathbf{1}_{A}$ $\mu$-almost surely, and $\mu$ is ergodic. ∎
@@
Let $P$ be a Markov kernel on a standard Borel space $(E,\mathcal{B})$. If $P$ is indecomposable, then $P$ admits at most one invariant probability measure. If one exists, it is ergodic.
@@
Assume by contradiction that $P$ admits two distinct invariant probability measures. By Proposition 5, there exist two mutually singular invariant probability measures $\mu_{1}$ and $\mu_{2}$. Hence there exists $A\in\mathcal{B}(E)$ such that
@@
Thus $B_{1}$ and $B_{2}$ are disjoint nontrivial absorbing measurable sets, contradicting the indecomposability of $P$. Uniqueness follows from Proposition 6. ∎
@@
On a standard Borel space, $\phi$–irreducibility in the sense of Meyn–Tweedie is a strictly stronger property than indecomposability, since $\phi$–irreducibility rules out the existence of nontrivial absorbing measurable sets. In general, the converse implication does not hold, as indecomposability alone does not preclude purely transient behavior. However, if an invariant probability measure exists, the structural decomposition theory of [9, Section 4.2] implies that indecomposability excludes the presence of more than one closed communicating class and therefore enforces irreducibility with respect to a maximal irreducibility measure $\psi$ in the sense of Meyn–Tweedie. In this sense, indecomposability and $\phi$–irreducibility become equivalent once the existence of an invariant probability measure is guaranteed. From this perspective, indecomposability identifies a purely structural obstruction to uniqueness of invariant probability measures, while stronger assumptions—such as the existence of small or petite sets or additional regularity of the transition kernel—are only required to address recurrence and ergodic convergence properties.
@@ 2.3 Stability via uniqueness
Theorem 16 (Stability induced by uniqueness).
Let $P$ be a quasi–Feller Markov kernel such that, for every $x\in E$, the empirical occupation measures $(\Lambda_{n})$ are $P_{x}$–almost surely tight. Under these assumptions, $P$ admits at least one invariant probability measure. If this invariant probability measure is unique, then the Markov chain $(X_{n})_{n\geq 0}$ is stable in the sense of Duflo.
@@ 4 Conclusion
We further show that, under irreducibility, indecomposability can be enforced through a resolvent–type kernel. The resolvent transforms finite–time reachability into a one–step positivity property at the level of invariant measures, forcing all invariant probability measures to charge a common $\sigma$–finite reference measure. Since distinct invariant ergodic measures are mutually singular, this excludes the possibility of multiple ergodic components and yields uniqueness without any appeal to return–time estimates or petite set constructions.